The Royal Fruit Company produces two types of fruit drinks. The first type is 65% pure fruit juice, and the second type is 90% pure fruit juice. The company is attempting to produce a fruit drink that contains 85% pure fruit juice. How many pints of each of the two existing types of drink must be used to make 80 pints of a mixture that is 85% pure fruit juice?
Answer by stanbon(75887)

The Royal Fruit Company produces two types of fruit drinks. The first type is 65% pure fruit juice, and the second type is 90% pure fruit juice. The company is attempting to produce a fruit drink that contains 85% pure fruit juice. How many pints of each of the two existing types of drink must be used to make 80 pints of a mixture that is 85% pure fruit juice?
-----------------------------
Equation:
pure + pure = pure
0.65x + 0.90(80-x) = 85*80
65x + 90*80 - 90x = 85*80
-25x = -5*80
x = (1/5)80
x = 16 pints (amt. of 65% solution needed)
----
80-x = 64 pints (amt. of 90% solution needed)
==========================
Cheers,
Stan H.